Let A = last nonzero digit of your ID number.

B = last 2 nonzero digits of your ID number. ID: 20-1511-267


1. A/An A-meter long metal pipe at B°C is heated until it increases its length by 0.0025 m at 102°C. (A) What is the initial temperature in Fahrenheit? (B) What is the initial temperature in kelvin?. (C) What is the coefficient of linear expansion of the metal?


2. Find the volume expansion of a brass block at B°C if it is heated to 100 °C. Its initial volume is 15 in3 . The coefficient of volume expansion for brass is 56 x 10-6/°C.


1
Expert's answer
2021-05-25T15:21:24-0400

A=7

B=67

1. (A) Initial temperature = B = 67o C

In Fahrenheit = "67\times\dfrac{9}{5}+32=152.6\degree F"

(B) In Kelvin = 67 + 273 = 340 K

(C) Initial length, "L=7\space m"

Change in length, "\Delta L=0.0025\space m"

Change in temperature, "\Delta T=102-67=35\degree C"

"\Delta L=L\alpha_L\Delta T"

"\alpha_L=\dfrac{\Delta L}{L\Delta T}=1.02\times10^{-5}\degree C^{-1}"


2.Initial temperature = 67o C

Final Temperature = 100o C

"\Delta T=100-67=33\degree C"

Initial Volume = 15 in3

"\alpha_V=56\times10^{-6}\degree C^{-1}"

"\Delta V=V\alpha_L\Delta T"

"\Delta V=0.02772\space in^3"